OpenStudy (anonymous):

The area of a parallelogram is 100 in.². Its height is 5 in. What are the lengths of the bases?

OpenStudy (anonymous):

Well the formula for the area of a parallelogram is Base * Height. (Proof @ http://en.wikipedia.org/wiki/Parallelogram#Area_formulas) Now Area is given and Height is given. Therefore you can easily find the base. 100=15*b; find b. :)

OpenStudy (anonymous):

so 15 is the answ

OpenStudy (anonymous):

Sorry. 100=5*b. so b=100/5. Therefore 20 is the answer.